For each of the following cost-output relationships, describe the shape (U-shape,

decreasing, increasing, constant) of the average total cost and marginal cost

functions (C = total cost, Q = output):

(a) C = 42,500,000 + 2550Q

(b) C = 8.48 + 0.65Q + .00220Q2


1
Expert's answer
2022-01-05T16:32:00-0500

(a) C = 42,500,000 + 2550Q,

ATC = C/Q = 42,500,000/Q + 2550, so it is decreasing,

MC = C'(Q) = 2550, so it is constant.


(b) "C = 8.48 + 0.65Q + 0.00220Q^2" ,

ATC = 8.48/Q + 0.65 + 0.0022Q, so it is U-shaped,

MC = 0.65 + 0.0044Q,so it is increasing.